OpenStudy (anonymous):

Yeah... i dont know how to factor polynomials. Factor. v^2 – 16

OpenStudy (anonymous):

v^2 - 16 can be written as v^2 - 4^2 The formula for a^2 - b^2 is (a+b)(a-b) So the expression can be factored as (v+4)(v-4)

OpenStudy (anonymous):

Did you know this formula?

OpenStudy (anonymous):

Oh I assume thats part of the fundementals of factoring polynomials, thanks!

OpenStudy (anonymous):

v^2-16 use perfect square rule which is a^2 - b^2 = (a+b)(a-b) =(v-4)(v+4) so when you expand it will become v^2 -4v + 4v -16.

OpenStudy (anonymous):

No I didnt.

OpenStudy (anonymous):

Hmm I see.

OpenStudy (anonymous):

You should memorise some basic formulae of factorization before solving questions.

OpenStudy (anonymous):

Will do.

OpenStudy (anonymous):

That's to check its right the -4v will be cancelled out by the +4v leaving v^2 - 16.

OpenStudy (anonymous):

oh thats cool , thanks guys!

OpenStudy (anonymous):

Your welcome :)